OpenStudy (anonymous):

A boat makes a 120-mile trip downstream in 3 hours but makes the return trip in 4 hours. What is the rate of the current? a)3 mph b)4 mph c)5 mph

OpenStudy (anonymous):

i would say c:) 120/2=60 60/3=20 and 60/4=15 20-15=5mph